# Stress Testing Generalization: How Minor Modifications Undermine Large Language Model Performance

### v1

@@ Abstract
This paper investigates the fragility of Large Language Models (LLMs) in generalizing to novel inputs, specifically focusing on minor perturbations in well-established benchmarks (e.g., slight changes in question format or distractor length). Despite high benchmark scores, LLMs exhibit significant accuracy drops and unexpected biases (e.g., preference for longer distractors) when faced with these minor but content-preserving modifications. For example, Qwen 2.5 1.5B’s MMLU score rises from 60 to 89 and drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when question types are changed, with a 6-point drop across all three modification categories. These analyses suggest that LLMs rely heavily on superficial cues rather than forming robust, abstract representations that generalize across formats, lexical variations, and irrelevant content shifts. This work aligns with the ACL 2025 theme track on the Generalization of NLP models, proposing a “Generalization Stress Test” to assess performance shifts under controlled perturbations. The study calls for reevaluating benchmarks and developing more reliable evaluation methodologies to capture LLM generalization abilities better.
Stress Testing Generalization: How Minor Modifications Undermine Large Language Model Performance
@@ 1 Introduction
While LLMs perform well in established benchmarks, concerns have been raised about the validity of these evaluations. Data contamination, where models unintentionally learn from benchmark data included in their training, can inflate performance estimates (Brown et al., 2020; Xu et al., 2024; Ravaut et al., 2024; Zhou et al., 2023). These issues suggest that patterns of existing benchmarks have been exposed. The existing benchmarks may not truly assess generalization.
Recent work has focused on uncovering the actual limits of LLM generalization. One direction involves the development of dynamic evaluation methods that modify the evaluation process on the fly (Zhu et al., 2024; Yu et al., 2024). Another approach emphasizes creating more challenging or adversarial test sets that push models beyond their current capabilities, such as MMLU-Pro (Wang et al., 2024) and GSM-Plus (Li et al., 2024a). A third line of inquiry involves introducing subtle modifications to benchmark datasets to test LLM robustness, such as altering the order of multiple-choice options or changing the format of questions (Zheng et al., 2024; Li et al., 2024b; Gupta et al., 2024; Alzahrani et al., 2024). While these approaches have contributed to a better understanding of LLM performance, they either increase the complexity of the evaluation or focus on relatively limited formatting changes like option ID adjustments.
@@
In this paper, we introduce a novel evaluation framework, Generalization Stress Tests, which examines LLM performance under three types of minor, content-preserving perturbations:
  • Changing question types (e.g., converting multiple-choice questions to boolean judgments).
@@
As shown in Figure 1, these simple modifications, surprisingly, lead to substantial performance degradation [^1] We test GSM-8K for noun replacement, as some MMLU cases lack irrelevant nouns.. We observe that LLMs struggle to generalize across varying option lengths, problem types, and noun replacements. For example, Qwen 2.5 1.5B’s MMLU score drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when question types are changed, with a 6-point drop across all three categories. These findings reveal a critical limitation: LLMs struggle to generalize beyond these shallow patterns and fail to replicate the human-like ability to ignore irrelevant format details.
@@ 2.2 Change Problem Type to Fairly Analyze LLMs’ Scoring Bias
Previous work found LLMs do not generalize to different option IDs in MCQ Zheng et al. (2024) and tried to solve this by changing the task to clozeAlzahrani et al. (2024). However, the cloze task reduces the expected value of selecting the correct answer. Therefore, we propose changing the multiple-choice questions to bool questions, requiring two judgments to be accurate, so that the difficulty of the questions is as similar as possible to that of multiple-choice questions.
As illustrated in Figure 3, We derive one true proposition that concludes with the right option and one false preposition that concludes with a randomly selected wrong option. LLMs need to judge both two propositions correctly.
@@ 3 Experiments
We perform evaluations on harness framework Gao et al. (2024) and adopt its default setting. We evaluate models of Llama3.1 series Dubey et al. (2024), Qwen2.5 series Yang et al. (2024b), and GPT4o. Llama3.1, and Qwen2.5 are the most powerful small models, while GPT4o is the most powerful LLM. We evaluate LLMs on MMLU Hendrycks et al. (2021), ARC-Challenge Clark et al. (2018), and GSM8k Cobbe et al. (2021). The first two are MCQ benchmarks, and the last consists of open-ended QAs. Refer to Appendix C for details.
@@ 3.1 Results of Altering Option Length
LLMs struggle to generalize across option length: From Table 1, it is evident that across all LLMs, from 1.5B to GPT4o, scores increase significantly when the length of the correct option is extended and decrease significantly when we make an incorrect option longer. Smaller models generalize worse.
Length matters, especially when we lengthen the right option. As shown in Table 2, changing the length can result in a difference of more than ten points in the RL setting.
@@
In Appendix D.1, LLMs tend to select the right option if we make all incorrect options longer.
@@ 4 Conclusion
This paper highlights the fragility of Large Language Models (LLMs) in generalizing to minor perturbations in benchmark tasks. Our experiments reveal that LLMs exhibit significant performance degradation when faced with slight changes in question format, option length, or irrelevant content shifts. These findings underscore that LLMs rely on superficial patterns rather than robust, generalizable reasoning. By introducing the "Generalization Stress Tests," we offer a novel framework for evaluating LLMs’ true generalization capabilities. This work aligns with the ACL 2025 theme on model generalization, advocating for developing more reliable benchmarks to assess LLMs beyond their superficial performance on traditional evaluation sets.
@@ A.1 Prompts
The prompt to control option length is: We concat the default prompt to one of the below prompts.
@@
We set the temperature to $0$.
@@ Appendix B Prompts in Replacing Irrelevant Nouns
We set temperature to $0.10.1$, top-p to $11$, top-k to $0$, and repetition_penalty to $0$.
@@ C.4 Budget
    We performed experiments with an H800 GPU; the total experiments cost about 1000 GPU hours.
@@ Appendix D Additional Results
D.1 Make ALL Wrong Options Longer.

### v2

# Large Language Models Badly Generalize across Option Length, Problem Types, and Irrelevant Noun Replacements

@@ Abstract
In this paper, we propose a “Generalization Stress Test” to assess Large Language Models’ (LLMs) generalization ability under slight and controlled perturbations, including option length, problem types, and irrelevant noun replacements. We achieve novel and significant findings that, despite high benchmark scores, LLMs exhibit severe accuracy drops and unexpected biases (e.g., preference for longer distractors) when faced with these minor but content-preserving modifications. For example, Qwen 2.5 1.5B’s MMLU score rises from 60 to 89 and drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when problem types are changed, with a 6-point drop across all three modification categories. These analyses suggest that LLMs rely heavily on superficial cues rather than forming robust, abstract representations that generalize across formats, lexical variations, and irrelevant content shifts.
Large Language Models Badly Generalize across Option Length, Problem Types, and Irrelevant Noun Replacements
@@ 1 Introduction
While LLMs perform well in established benchmarks, concerns have been raised about the validity of these evaluations Chen et al. (2023); Ye et al. (2023). Data contamination, where models unintentionally learn from benchmark data included in their training, can inflate performance estimates (Brown et al., 2020; Xu et al., 2024; Ravaut et al., 2024; Zhou et al., 2023). These issues suggest that existing benchmarks have exposed patterns and may not truly assess generalization.
Recent work has focused on uncovering the actual limits of LLM generalization. One direction involves the development of dynamic evaluation methods that modify the evaluation process on the fly (Zhu et al., 2024; Yu et al., 2024). Another approach emphasizes creating more challenging or adversarial test sets that push models beyond their current capabilities, such as MMLU-Pro (Wang et al., 2024) and GSM-Plus (Li et al., 2024a). A third line of inquiry involves introducing subtle modifications to benchmark datasets to test LLM robustness, such as altering the order of multiple-choice options or changing the format of questions (Zheng et al., 2024; Li et al., 2024b; Gupta et al., 2024; Alzahrani et al., 2024). While these approaches have contributed to a better understanding of LLM performance, they either totally change the original problems, increase the complexity of the evaluation, or focus on relatively limited formatting changes like option ID adjustments.
@@
We find serious biases of recent SoTa LLMs to common patterns by introducing an evaluation framework, Generalization Stress Tests, which examines LLMs under three types of minor, content-preserving perturbations:
@@
  • Changing problem types (e.g., converting multiple-choice questions to boolean questions).
@@
As shown in Figure 1, these simple modifications, surprisingly, lead to substantial performance degradation [^1] We test GSM-8K for noun replacement, as some MMLU cases lack irrelevant nouns.. We observe that LLMs struggle to generalize across varying option lengths, problem types, and noun replacements. For example, Qwen 2.5 1.5B’s MMLU score drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when question types are changed, with a 6-point drop across all three categories. These findings reveal a critical limitation: LLMs are biased to specific irrelevant patterns and fail to replicate the human-like ability to ignore irrelevant format details.
@@ 2.2 Change Problem Type to Fairly Analyze LLMs’ Scoring Bias
Previous work found LLMs do not generalize to different option IDs in MCQ Zheng et al. (2024) and tried to solve this by changing the task to cloze Alzahrani et al. (2024). However, the cloze task reduces the expected value of selecting the correct answer. Therefore, we propose changing the multiple-choice questions to Boolean questions, requiring two judgments to be accurate, so that the difficulty of the questions is as similar as possible to that of multiple-choice questions.
As illustrated in Figure 3, we derive one true proposition that concludes with the right option and one false proposition that is a randomly selected wrong option.
@@ 3 Experiments
We perform evaluations on harness framework Gao et al. (2024) and adopt its default setting. We evaluate models of Llama3.1 series Dubey et al. (2024), Qwen2.5 series Yang et al. (2024b), and GPT4o. Llama3.1, and Qwen2.5 are the most powerful small models, while GPT4o is the most powerful LLM. We evaluate LLMs on MMLU Hendrycks et al. (2021), ARC-Challenge Clark et al. (2018), and GSM8k Cobbe et al. (2021). The first two are MCQ benchmarks, and the last consists of open-ended QAs. Refer to Appendix C for detailed experimental setups.
@@ 3.1 Results of Altering Option Length
LLMs struggle to generalize across option length: From Table 1, it is evident that across all LLMs, from 1.5B to GPT4o, scores increase significantly when the length of the correct option is extended and decrease significantly when we make an incorrect option longer. Smaller models generalize even worse. In Appendix D.1, we introduce another setting of making all options longer, in which our finding that LLMs are biased towards the longer option persists.
Length matters, especially when we lengthen the right option. As shown in Table 2, changing the length can result in a difference of more than 10 points in the RL setting.
@@
Another intriguing finding is that LLMs tend to select the right option if we make all incorrect options longer, refer to Appendix D.2.
@@ 3.3 Results of Replacing Irrelevant Nouns
3.4 Reasoning Behind Accuracy Drops
The above analysis reveals that LLMs are severely biased to common but irrelevant patterns.
@@ 4 Conclusion
This paper finds that LLMs exhibit significant performance degradation when faced with slight changes in question format, option length, or irrelevant content shifts. These findings underscore that LLMs rely on superficial patterns rather than robust, generalizable reasoning. By introducing the "Generalization Stress Tests," we offer novel ideas towards evaluating LLMs’ true generalization capabilities.
@@ A.1 Prompts
The prompt to control option length is: We concatenate the default prompt to one of the following prompts.
@@
We set the temperature to 0, and the other setting is the same as the default.
@@ Appendix B Prompts in Replacing Irrelevant Nouns
We set temperature to $0.1$, top-p to $1$, top-k to 0, and repetition_penalty to 0.
@@ C.4 Budget
    We performed experiments with an H800 GPU; the total cost of the experiments was about 1000 GPU hours.
@@ Appendix D Additional Results
D.1 Making All Options longer
Benchmark Model Origin AL RL WL
MMLU Qwen2.5 1.5B 60.3 54.7 89.0 36.3
Qwen2.5 7B 73.7 69.2 90.1 55.6
Qwen2.5 72B 85.4 81.3 94.1 75.6
LLaMa3.1 8B 65.5 64.3 85.6 53.6
LLaMa3.1 70B 78.8 76.0 93.6 70.6
ARC-C Qwen2.5 1.5B 77.3 67.3 88.9 68.1
Qwen2.5 7B 90.0 85.3 94.3 84.0
Qwen2.5 72B 95.8 93.1 97.2 94.4
LLaMa3.1 8B 78.1 78.6 85.2 74.7
LLaMa3.1 70B 91.8 89.9 96.3 90.8
Table 6: Performance on altering option length. AL refers to lengthening all options. RL refers to lengthening the right option. WL refers to lengthening the wrong option. The values are percentages.
We can see from Table 6 that LLaMa is more robust than Qwen, and larger models are more robust than smaller models, when we make all options longer. Besides, even if we introduce the setting of AL, our conclusion that LLMs are vulnerable to option lengths and biased to long options is not changed.
D.2 Make All Wrong Options Longer